OpenStudy (anonymous):

Use the 4 step method to find the inverse of f-1(x) ----f(x)= 2x-1. Please show work

OpenStudy (anonymous):

f(x)= 2x-1 y = 2x-1 y+1=2x x= (y+1)/2 f-1(x)=(x+1)/2 [x replaced by y]

OpenStudy (anonymous):

yes but we started with f-1 didn't we? So the answer we get shouldn't be f-1 right? should just be f(x)?

OpenStudy (anonymous):

the given is f(x)=2x+1, the question is "find and simplfy" and the prompt is f-1(x) (hint:4-step method)

OpenStudy (anonymous):

yes do that f(x)= 2x-1 f(x)+1=2x x= (f(x)+1)/2 f-1(x)=(x+1)/2